OpenStudy (anonymous):

James opened a money-market account that paid 5% simple interest yearly. He started the account with $7,000 and made no further deposits. When he closed the account, he had earned $1,750 in interest. How long did he keep his account open?

OpenStudy (anonymous):

i =PRT T=i/PR T= 1750/(7000*0.05) = 5 years